An experienced bricklayer can construct a small wall in 4 hours. An apprentice can complete the job in 6 hours. Find how long it takes if they work together.
Answer by ptaylor(2198)

Let x=amount of time it takes if they both work together
Together they work at the rate of 1/x wall per hour
Experienced bricklayer works at the rate of 1/4 wall per hour
Apprentice works at the rate of 1/6 wall per hour
So our equation to solve is:
1/4 + 1/6= 1/x multiply each term by 12x
3x+2x=12
5x=12
x=2.4 hours---time it takes if they both work together
Check
In 2.4 hours, experienced bricklayer lays (1/4)*2.4=0.6 of a wall
In 2.4 hours apprentice lays (1/6)*2.4=0.4 of a wall
0.6+0.4=1 (wall, that is)
Hope this helps--ptaylor